APPRENTICE

Name (18)  @uchicago.edu
Padraic James Bartlett  padraic
Ryan Dietzman  gammon
Nabeel Ebeid  nabeel
Matthew Gonwa  gonwa
Molly Hales  mkhales
Milan Kidd  mkidd
Yee Fay Lim*  yflim
Neil Lutz  njlutz
James Marrone  jmar
Jonathan Nieder  jrnieder
Pawan Patel  pawan
Edward T. Ruggeri  ruggeri
Jenya Sapir  jmsapir
Matthew Tunnell  mtunnell
Walter Ulmer  ulmer
Maya Vinokour  vinokour
Muhammad Waliji  mhwaliji
Jonathan Williams  jonathan

*Sit in Sit in (8 week)

Kevin Choi kylchoi

Helping Paul Sally (8 week)

Jeff Kwan hell-fire-wall@yahoo.com

YSP

Name (21)  Program @uchicago.edu
Sam Altschul  9-10  salt
Andrea Byrnes  7-8  aebyrnes
Amanda Comi  9-10  abc
Milind Desai  7-8  mdesai
Alexander Frankel  11-12  afrankel
Igor Gorodezky  11-12  igor2006
Ilya Grigoriev  9-10  ilyag
Moshe Hoffman  7-8  mozes
Martin Jankowiak  9-10  martinj
Evan Jenkins  11-12  ejenkins
Phillip Kalikman  9-10  plk
Brian Keto  7-8  bketo
Thomas Koberda  11-12  koberdat
Deniz Sehime Kusefoglu  7-8  deniz
Sara Lapan  9-10  slapan
Jeff Meyer  7-8  jsmeyer
Kyle Ormsby  11-12  kmo
James Pascaleff  11-12  jpascale
Allie Shapiro  7-8  allie
Weiwei Tan  9-10  cora7
Fang Zhao  7-8  roseanne

SESAME

Name (25)  @uchicago.edu
Philip Ascher  pa7
William Cavendish  wpcavend
Trevor Davis  trevorld
Daniel Arthur Dickinson  dand
Jesse Farmer  farmerje
Stephen David Flood  sflood
William T. Gignac  wtgignac
Shawn Henry  shawnh
Sayer Herin  sayerh
Josh Jones  jonesj
David Levinson  levinson
Paul C. Mantz  pcmantz
Jonathon Mason  jmason
Duff Neill  duff
Joe Ochiltree  jjo
Annie Roberts  redoak84
Rodrigo Rodriguez  rrodrigo
Jeff Schonert  jeffrey1
Jason Shapiro  shapiro
Rafal Sobota  rsobota
Sonya Swanson  sonja
Geoffrey Thompson  gzt
Alex V. Tomasik  tomasik
Sokol Vako  vako
Calder Wishne  calder